OpenStudy (anonymous):

Margaret drove to a business appointment at 50 mph. Her average speed on the return trip was 40 mph. The return trip took 1/4 hr longer because of heavy traffic. How far did she travel to the appointment?

OpenStudy (anonymous):

Let x be the distance to the business appointment.\[r* t=d,t=\frac{d}{r} \]Solve the following for x:\[\frac{x}{40}-\frac{x}{50}=\frac{1}{4} \]x = 50 miles